OpenStudy (anonymous):

what is the derivative of 1/t^2

OpenStudy (anonymous):

with respect to what??

OpenStudy (anonymous):

d/dx(1/t^2) = \[-\frac{2}{t^3}\]

OpenStudy (anonymous):

implied t argument

OpenStudy (anonymous):

with respect to y.

OpenStudy (anonymous):

so the ans will be -2/t^2*(dt/dy)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

y = 1/t^2 + t^1/2 this is the real equation. i got the second part but i cant get the fist part

OpenStudy (anonymous):

actually first of all y find out d/dt(1/t^2) as there is no terms of y so to make it w.r.t y just multiply wit dy/dy..then actual equ stands d/dy(1/t^2)*(dy/dt).

OpenStudy (anonymous):

the ans will be y=-2/t^3*(dy/dt)+1/2*(dy/dt)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

most welcome!!

OpenStudy (anonymous):

i miss some terms in 2nd part ur ans is y=-2/t^3*(dy/dt)+1/2*1/t^(1/2)*(dy/dt)